OpenStudy (anonymous):

what is the simplified form of the product x^2+8x+15/x-4 times x^2 -16/2x+6

OpenStudy (anonymous):

@autumnlovescorbin

OpenStudy (anonymous):

Are there any answers to choose from?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Can you post them?

OpenStudy (anonymous):

(x-5)(x+4)/2 (x+5)(x+4)/2 (x+5)(x-4)/2 (x-5)(x-4)/2

OpenStudy (anonymous):

u are there

OpenStudy (anonymous):

are you sure the 4 is not a 3 in the answers?

OpenStudy (anonymous):

yeah its 4

OpenStudy (anonymous):

k just a sec

OpenStudy (anonymous):

are u still there

OpenStudy (anonymous):

I am stumped I am trying to get more help

OpenStudy (anonymous):

I have to go, I will leave this for others Im sorry I was of no help

OpenStudy (anonymous):

Hello there

OpenStudy (anonymous):

him

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

can u helpp me

OpenStudy (anonymous):

\[x ^{2}+8x+15 / x-4 \times x ^{2} -16 /2x+6\]

OpenStudy (anonymous):

ok so if im correct this is what you had

OpenStudy (anonymous):

Trinomial: x^2 + 8x + 15 = (x + 3)(x + 5) (x^2 - 16) - (x - 4)(x + 4) Finally, after simplification by ((x - 4) and (x + 3) (x + 5)(x + 4)/2

OpenStudy (anonymous):

thank u